OpenStudy (anonymous):

Mike has 3 more dimes than nickels, and 5 quarters fewer than twice the number of nickels. In total he has $3.60. How many quarters does he have?

OpenStudy (anonymous):

3(10x) + 5x +5(25x)-20x=3600 x=25 1 quarter

OpenStudy (anonymous):

i am not sure about the answer though

OpenStudy (anonymous):

Wouldn't it be 10(x + 3)...etc, cause it says 3 MORE dimes than nickels not sure to add or multiply

OpenStudy (anonymous):

Let x=amount of nickles Then amount of dimes=(x+3) And amount of quarters=(2x-5) We then have 5(x)+10(x+3)+25(2x-5)=360 Solve: 65x=455 x=7 (we have seven nickles) So we have 9 (2*7-5) = 9 quarters

OpenStudy (anonymous):

Thanks bro !!! Your the best man!!